q:an automobile moving with a velocity of 54 km/hr is brought to rest by a distance of 9 m by the application of brakes.find the retardation assuming it to be uniform.

Grade:11

1 Answers

Sanjeev Malik IIT BHU
12 years ago

use the formula 2as= v2-u2

where a= acceleration or retardation   s=distance    v=final speed   u = initial speed

final speed =0 as automobile comes to rest

as distance is in metres , so also take speed in m/s.

u= 54km/h=54*1000/3600=15m/s

applying formula,

2*a*9=0-152

so a= 225/18=12.5 m/s2

 

 

approve by clicking 'yes" below..plz